OpenStudy (oladimeji):

An ideal gas is in equilibrium at initial state with temperature T=137oC, pressure P = 0.75Pa and volume V = 0.75 m3. If there is a change in state in which the gas undergoes an isothermal process to a final state of equilibrium during which the volume is doubled. Calculate the temperature and pressure of the gas at this final state.

OpenStudy (anonymous):

since the gas is in an equilibrium state therefore we can consider that Both Temp.& Pressure hav constant values so we can apply both Boyle's & Charle's law Boyle's law: V\[\alpha\] 1/P At constant T so the pressure will be halved charle's law: V\[\alpha\] T At constant P so the Temp. will be doubled Too

OpenStudy (oladimeji):

thanks but its not working. i tried \[V_{1}P_{1}\div T_{1} = V_{2}P_{2}\div T_{2}\] but its not working. How can i get both pressure and temperature at different times when they are unknown in the same equation.

OpenStudy (anonymous):

well u hav the volume & the volume has an equation with pressure & other with Temp. each equation is single with itself why u mixed them ?!!

OpenStudy (oladimeji):

Please express the equation i should use, if i can't use the ideal gas situation which is not required or better still can u solve it for me and show me the steps so i can learn?

OpenStudy (oladimeji):

kindly reply? i will be back online in some minutes back

OpenStudy (anonymous):

\[v1\div v2=p2\div p1\] & \[v1\div v2=t1\div t2\]

OpenStudy (anonymous):

Given that the reaction is isothermal we know that we have no variations in temperature, so T=T2=137C.Now when the pressure doubles we set P2=2P=1.5Pa.From boyle's law now we have P2*V2=P*V (the boyles law states that in an ideal gas the product of pressure and volume is constant) from the above we get V2=0.375Pa

OpenStudy (ujjwal):

the initial and final temperature is same since it is an isothermal process.. i.e. T1=T2 .. Now use P1/V1 = P2/V2

OpenStudy (anonymous):

We can surely use the ideal gas law here. First, let's fix our states as best we can with the information given in the problem. 1) Isothermal process: \(T_1 = T_2\). 2) Volume is doubled: \(V_1 = {1 \over 2}V_2\) 3) Assume a closed system: \(n_1 = n_2\). Therefore, from the ideal gas law, we get\[{P_1 V_1 \over n_1 R T_1} = {P_2 V_2 \over n_2 R T_2}\]From #1 and #3, we can simplify the above expression to be\[{P_1 V_1 \over T_1} = {P_2 V_2 \over T_1} \rightarrow P_1 V_1 = P_2V_2\]From #2, we can further simplify it to be\[{P_1 V_1} = {P_2 (2V_1) }\]\[P_2 = {P_1 V_1 \over 2V_1} = {P_1 \over 2}\]The final expression is Boyle's Law, but it's nice to be able to work through it from the ideal gas law in-case we forget under what conditions Boyle's Law holds (isothermal and constant amount).

OpenStudy (oladimeji):

thanks so far